OpenStudy (anonymous):

find the difference of x squared -2 over x-2 and 2x over x-2

OpenStudy (anonymous):

\[\frac{x^2-2}{x-2} - \frac{2x}{x-2}\]?

OpenStudy (anonymous):

if so , then : \[=\frac{x^2-2x-2}{x-2} = \frac{(x-2)(x+1)}{x-2}\] \[= x+1\] ^_^

OpenStudy (anonymous):

factoring is not right :p

OpenStudy (anonymous):

(x^2 -2x-2)/ (x-2) is as far as you can go. pretty small and disappointing question , usually with questions like those they have lots of factors that cancel and at the end you are left with like one term as you answer

OpenStudy (anonymous):

oh oh oh, riiiight lol >_<! I'm sorry

OpenStudy (anonymous):

thank you ^_^

OpenStudy (anonymous):

pammie dear, my answer is incorrect